# WeAudit: Scaffolding User Auditors and AI Practitioners in Auditing Generative AI

### v3

@@ Abstract.
copyright: none
@@ 1. INTRODUCTION
End users often surface harmful behaviors in Artificial Intelligence (AI) systems that are otherwise overlooked by AI practitioners (Shen et al., 2021; DeVos et al., 2022; Lam et al., 2022; Mack et al., 2024; Shelby et al., 2024; Li et al., 2023; Deng et al., 2023a; Solyst et al., 2023; Kieslich et al., 2024; Mun et al., 2024; Maldaner et al., 2024). Past research shows that AI users can detect and raise awareness of biased and harmful AI behaviors by leveraging their lived experiences, building on one anothers’ findings, and generating and testing hypotheses about AI harms (Shen et al., 2021; DeVos et al., 2022). There has been rapidly growing interest among both researchers and practitioners in engaging end users in auditing AI products and services: one line of recent research has explored systems and processes to enable user engagement in auditing (Attenberg et al., 2015; Cabrera et al., 2021; Kiela et al., 2021; Nushi et al., 2018; Ochigame and Ye, 2021; Lam et al., 2022; Deng et al., 2023b), while another has focused on understanding industry AI practitioners’ needs and challenges around effective user engagement (Deng et al., 2023a; Ojewale et al., 2024; Deng et al., 2024). However, a critical gap remains in connecting the needs, challenges, and perspectives of these two groups of stakeholders. How might we develop tools and processes to effectively scaffold end-user engagement in AI audits, while ensuring their findings are useful and actionable for AI practitioners?
@@
To understand both end users’ and AI practitioners’ needs for support in user auditing, we first conducted a formative study with 11 end users, and seven industry AI practitioners with prior experience engaging end users in AI auditing. These formative studies led to the development of a set of design goals, and a corresponding WeAudit workflow. We then instantiated the WeAudit workflow through a system that supports and scaffolds users in investigating, deliberating, and reporting perceived harms and biases in text-to-image (T2I) AI systems. To understand the strengths and limitations of WeAudit in supporting user auditing, we first conducted a three-week user study with 45 users to audit Stable Diffusion, an open source T2I system, and reflect on their experience using the tool. We then conducted semi-structured interviews with 10 industry GenAI practitioners to evaluate the design of WeAudit, grounded in data and user audit reports from our user study.
@@ 2.1. Engaging End Users in Auditing AI Systems for Harmful Behaviors
While most AI audits are still conducted by experts such as researchers (Buolamwini and Gebru, 2018; Kaufmann et al., 2011; Cramer et al., 2018; Eslami et al., 2017, 2019; Raji et al., 2020) and AI practitioners (Sweeney, 2013; Sandvig et al., 2014; Metaxa et al., 2021; Chowdhury and Williams, 2021), end users can often identify and raise awareness of harmful AI behaviors in systems that impact their lives, complementing expert-led AI audit efforts (Shen et al., 2021; DeVos et al., 2022). For example, end-users often discover harmful biases in Text to Image (T2I) generative AI systems that expert auditors fail to detect (Mack et al., 2024; Zhang et al., 2024; Shelby et al., 2024; Mim et al., 2024). When Mack et al. engaged people with disabilities to review and reflect on images generated by T2I systems they uncovered several societal stereotypes, such as “perpetuating broader narratives in society around disabled people as primarily using wheelchairs, being sad and lonely, incapable, and inactive” (Mack et al., 2024).
@@
Recognizing the power of users in AI auditing, researchers in CSCW, HCI, and AI have begun exploring tools and processes to support more user-engaged approaches to AI auditing, where users actively participate in the auditing process alongside efforts from AI practitioners or other technical experts (Shen et al., 2021; DeVos et al., 2022; Lam et al., 2022; Attenberg et al., 2015; Cabrera et al., 2021; Kiela et al., 2021; Nushi et al., 2018; Kieslich et al., 2024). For example, DeVrio et al. conducted behavioral studies to investigate how users, both individually and collectively, are able to surface harmful algorithmic behaviors that formal or expert-led audits may miss (DeVos et al., 2022). Lam et al. developed an “End-user Audits” framework, instantiated in a system called IndieLabel, to engage non-technical users in auditing sentiment analysis classification AI models (Lam et al., 2022).
@@
In parallel, several major technology companies have begun experimenting with user-engaged auditing approaches to identify harmful and biased outputs in their AI products and services (Pistilli, 2022; OpenAI, 2022a; Warkentin and Woodward, 2022; Chowdhury and Williams, 2021; Kiela et al., 2021). In 2021, Twitter launched its first “algorithmic bias bounty” challenge, inviting users to identify harmful biases in its image-cropping algorithm (Chowdhury and Williams, 2021). Following this, OpenAI introduced a “Feedback Contest” encouraging users to provide feedback on problematic model outputs during their interactions with ChatGPT. Many corporations with generative AI products have also turned to “red-teaming” events, where users rigorously test models to identify vulnerabilities and harmful biases (such as the AI village at DEF CON, a hacking conference) (CON, 2024; McGregor et al., 2024). In response to calls from government agencies like the Executive Order on Safe, Secure, and Trustworthy Development and Use of AI (The White House, 2023), many civil society organizations has called for technology companies to engage external domain experts, including end users, in auditing their generative AI products (Groves et al., 2023; Bogen and Winecoff, [n. d.]; Park, [n. d.]; Feffer et al., 2024).
@@
A small but growing body of research has begun exploring the needs and challenges faced by AI practitioners in conducting user-engaged AI audits. For example, through interviews with 35 AI audit practitioners, Ojewale et al. identified the need for tools to support AI practitioners in engaging those directly impacted by AI deployments, offering insights into how to better design participatory AI tools and processes with accountability (Ojewale et al., 2024). In a study with industry AI practitioners who are engaging users in auditing their AI systems, Deng et al. found that practitioners often repurpose crowdsourcing platforms for user audits but struggle to scaffold users toward productive auditing strategies and to derive actionable insights from users’ audit outcomes (Deng et al., 2023a).
@@ 3.1.1. Think-aloud study with end users
To understand the challenges end users encounter when participating in GenAI audits, and to identify opportunities to better support them, we first conducted think-alouds with 11 end users. For these formative studies, we aimed to recruit participants spanning diverse demographics and levels of technical literacy. To this end, we recruited participants through Craigslist and Nextdoor. 10 of our 11 participants had not previously used any T2I systesm; participants ranged from 18 years old to 80 years old, with 47.45 average. 2 participants has high school degrees, 2 Associate degrees, 5 bachelor degrees, 1 master degree, and 1 Ph.D. We include detailed participants descriptions in Table 3 under the Appendix A.1.
@@
After this initial phase, for the next 20 minutes participants continued their testing, but now with access to a set of three low-fidelity prototypes intended to support their process. The first prototype was a pairwise comparison feature, simulated by placing two browser windows side-by-side. We hypothesized that viewing prompt outputs side-by-side would help participants, especially those unfamiliar with T2I systems, understand how prompt changes affect the output. The second prototype was a sidebar displaying a sample user audit report in text format. This report featured an election-themed prompt and image pair, with a brief sentence explaining why the AI-generated image could be harmful. We hypothesized that this example could inspire participants to search for similar cases and better articulate the harms they identified. The third prototype was a T2I harm taxonomy, mapping Shelby et al.’s sociotechnical AI harm taxonomy (Shelby et al., 2023) to concrete examples of harmful T2I behaviors, based on examples from (Luccioni et al., 2023). We hypothesized that this taxonomy would help users brainstorm a broader range of potential harms to various communities and use terminology to describe these harms effectively. We include screenshots of each prototype in Figure 3.
@@ 3.1.2. Semi-structured Interview with Industry GenAI Practitioners
To understand industry GenAI developers’ needs for support in effectively engaging end-users in AI auditing, we conducted semi-structured interviews with 7 industry AI developers. To recruit participants, we adopted a purposive sampling approach (Campbell et al., 2020), with the aim of recruiting industry practitioners who had already attempted to engage end users in evaluating and auditing their GenAI products  [^2] Practitioners frame this activity differently depending on their organization, in some cases using terminology like “red-teaming”.. We recruited participants through social media (e.g., LinkedIn and X), and via direct contacts at large technology companies. Overall, industry practitioners come from diverse backgrounds, working in technical, user-facing, and managerial roles on various types of AI products powered by Large Language Models (LLMs) or other types of generative AI. Notably, all of them are currently using crowdsourcing platforms to engage individuals outside their AI teams in auditing or red-teaming their AI products and services. Table 5 in the Appendix A.1 provides an overview of participants’ backgrounds  [^3] Following prior work studying responsible AI practices in industry (Madaio et al., 2020; Rakova et al., 2021; Deng et al., 2023c), to protect participants’ confidentiality, we omit potentially identifying demographic details such as gender and age, and abstract our descriptions of participants’ companies and roles..
@@ 4.1. The WeAudit Workflow
As shown in Figure 4, the WeAudit workflow is organized into two intersecting, iterative loops: Investigate and Deliberate. The overall loop structure of this workflow is informed by existing models of information foraging and sensemaking (Pirolli and Card, 1999), and these loops are further broken down into six sub-activities: Explore, Inspect, Reflect, Report, Discuss, and Verify. In the Investigate loop, user auditors may begin investigating AI harms by exploring prompts that could lead to harmful outputs (DG1, DG3). Auditors inspect the resulting AI-generated outputs (DG1), and reflect on potential biases or harmful impacts, drawing upon their unique knowledge and experiences (DG2). Auditors can then report their findings, documenting potential harms they identified (DG4). In the Deliberate loop, user auditors can discuss their findings with other auditors or the broader public (DG5). This collaboration allows for a more holistic understanding of AI-related harms, leveraging user auditors’ diverse perspectives. Additionally, reported findings can be verified by other auditors or the public (DG6), to provide a sense of whether a given observation is perceived as potential harmful by multiple people. In Table 1 above, we further illustrate how each component of the WeAudit workflow reflects the Design Goals.
Design Goals How the WeAudit Workflow Reflects the Design Goals
DG1: Hypothesis generation and validation The overall Investigate loop aligns with this design goal by supporting users in testing different prompts and identifying potential harms.
DG2: Incorporation of personal experience The Reflect activity that cuts across Investigate and Deliberate aims to encourage users to bring in their personal experiences and identities.
DG3: Efficient and creative audits The Explore step aims to support users to test different prompts, while avoiding over-reliance on existing audit examples.
DG4: Communication of harms The Report step aims at helping users clearly communicate perceived AI harms, to support discussion and collective sensemaking among user auditors, and to enable developers to extract actionable insights.
DG5: Discussion and deliberation The Discuss and Verify activities aim to promote user collaboration and deliberation.
DG6: Verification of reports The Verify step aims to support verification of audit reports, while being cautious not to disempower minority voices.
Table 1. How loops and sub-activities in the WeAudit workflow reflect the Design Goals in Section 3.2
In the workflow, we use the circular arrows to highlight that the entire workflow is iterative, following prior sensemaking research (Pirolli and Card, 1999). For example, new insights that arise during discussion or verification may inspire user auditors to conduct new investigations. Moreover, this iterative workflow engages user auditors in fluidly moving back-and-forth between individual and collective activities, where individual efforts contribute to the collective understanding of AI behaviors and vice versa. Throughout the entire process, user auditors are encouraged to reflect on potential harms based on their own lived experiences and identities. Since the “reflect” and “report” activities function as bridges between Investigate and Deliberate loops, we place them within the intersection of these loops.
@@ 4.2. The WeAudit System
Figure 5. Interface for features: (a) Pairwise Comparison (Section 4.2.1), (b) Prompt History Sidebar (Section 4.2.2), (c) Worked Examples Repository (Section 4.2.3), and (d) Social Augmentation (Section 4.2.4)
@@ 4.2.4. Social Augmentation
In formative study, we collected initial evidences on participants found “sidebar examples” from other user auditors being useful (see Section 3). Prior research in CSCW, crowdsourcing, and sensemaking also shows that reviewing other users’ behaviors can affect users motivations and strategies for completing a task (Morris and Horvitz, 2007; Amershi and Morris, 2008; Dow et al., 2012). Inspired by this, we provide users with an aggregated view of what other users have been reporting. User auditors can click on the ”What are other users auditing?” button to view a distribution of “affected groups” (e.g., religion or sexual orientation) and “types of harms” (e.g., stereotyping social groups or economic loss) currently covered by submitted audit reports, with ones that are “underexplored” and ones that are “most explored” highlighted (Figure 5). By clicking each tag, users can view the specific user audit posts associated with those tags. Through this form of social augmentation, our goal is to encourage users to explore auditing directions that have been less frequently explored by others (DG3), drawing upon their own unique identities, experiences, and knowledge (DG2).
Figure 6. Interface for features: (e) Audit Report Portal (Section 4.2.5), an example of audit report authored by a user auditor in our user study through the Audit Report Portal, and (f) Audit Discussion Forum (Section 4.2.6). Please refer to Section 4.2.5 to review the concrete questions in the Audit Report Portal.

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. INTRODUCTION
End users often surface harmful behaviors in Artificial Intelligence (AI) systems that are otherwise overlooked by AI practitioners (Shen et al., 2021b; DeVos et al., 2022; Lam et al., 2022; Mack et al., 2024; Shelby et al., 2024; Li et al., 2023; Deng et al., 2023a; Solyst et al., 2023; Kieslich et al., 2024; Mun et al., 2024). Past research shows that AI users can detect and raise awareness of biased and harmful AI behaviors by leveraging their lived experiences, building on one anothers’ findings, and generating and testing hypotheses about AI harms (Shen et al., 2021b; DeVos et al., 2022). There has been rapidly growing interest among both researchers and practitioners in engaging end users in auditing AI products and services: one line of recent research has explored systems and processes to enable user engagement in auditing (Attenberg et al., 2015; Cabrera et al., 2021; Kiela et al., 2021; Nushi et al., 2018; Ochigame and Ye, 2021; Lam et al., 2022; Maldaner et al., 2024; Claire et al., 2024), while another has focused on understanding industry AI practitioners’ needs and challenges around effective user engagement (Deng et al., 2023a; Ojewale et al., 2024; Wang et al., 2023; Madaio et al., 2024, 2021; Deng et al., 2024b, a). However, a critical gap remains in connecting the needs, challenges, and perspectives of these two groups of stakeholders. How might we develop tools and processes to effectively scaffold end-user engagement in AI audits, while ensuring their findings are useful and actionable for AI practitioners?
@@
To understand both end users’ and AI practitioners’ needs for support in user auditing, we first conducted a formative study with 11 end users (U01-U11), and seven industry AI practitioners with prior experience engaging end users in AI auditing (P01-P07). These formative studies led to the development of a set of design goals, and a corresponding WeAudit workflow that draws upon ideas from CSCW research on crowdsourcing and sensemaking (Kittur et al., 2013; Cabrera et al., 2022; Pirolli and Card, 1999; Chan et al., 2018). We then instantiated the WeAudit workflow through a system that supports and scaffolds users in investigating, deliberating, and reporting perceived harms and biases in text-to-image (T2I) AI systems. To understand the strengths and limitations of WeAudit in supporting user auditing, we first conducted a three-week user study with 45 users (F01-F45) to audit Stable Diffusion, an open source T2I system, and reflect on their experience using the tool. We then conducted semi-structured interviews with 10 industry GenAI practitioners (a subset of P01-P14) to evaluate the design of WeAudit, grounded in data and user audit reports from our user study.
@@ 2.1. Engaging End Users in Auditing AI Systems for Harmful Behaviors
While most AI audits are still conducted by experts such as researchers (Buolamwini and Gebru, 2018; Kaufmann et al., 2011; Cramer et al., 2018; Eslami et al., 2017, 2019; Raji et al., 2020) and AI practitioners (Sweeney, 2013; Sandvig et al., 2014; Metaxa et al., 2021; Chowdhury and Williams, 2021), end users can often identify and raise awareness of harmful AI behaviors in systems that impact their lives, complementing expert-led AI audit efforts (Shen et al., 2021b; DeVos et al., 2022). For example, end-users often discover harmful biases in Text to Image (T2I) generative AI systems that expert auditors fail to detect (Mack et al., 2024; Zhang et al., 2024; Shelby et al., 2024; Mim et al., 2024). When Mack et al. engaged people with disabilities to review and reflect on images generated by T2I systems they uncovered several societal stereotypes, such as “perpetuating broader narratives in society around disabled people as primarily using wheelchairs, being sad and lonely, incapable, and inactive” (Mack et al., 2024).
@@
Recognizing the power of users in AI auditing, researchers in CSCW, HCI, and AI have begun exploring tools and processes to support more user-engaged approaches to AI auditing, where users actively participate in the auditing process alongside efforts from AI practitioners or other technical experts (Shen et al., 2021b; DeVos et al., 2022; Lam et al., 2022; Attenberg et al., 2015; Cabrera et al., 2021; Kiela et al., 2021; Nushi et al., 2018; Kieslich et al., 2024). For example, DeVrio et al. conducted think-alouds, diary studies, and workshops to investigate how users, both individually and collectively, are able to surface harmful algorithmic behaviors that formal or expert-led audits may miss (DeVos et al., 2022). Lam et al. developed an “End-user Audits” framework, instantiated in a system called IndieLabel, to engage non-technical users in auditing sentiment analysis classification AI models (Lam et al., 2022).
@@
In parallel, several major technology companies have begun experimenting with user-engaged auditing approaches to identify harmful and biased outputs in their AI products and services (Pistilli, 2022; OpenAI, 2022a; Warkentin and Woodward, 2022; Chowdhury and Williams, 2021; Kiela et al., 2021). In 2021, Twitter launched its first “algorithmic bias bounty” challenge, inviting users to identify harmful biases in its image-cropping algorithm (Chowdhury and Williams, 2021). Following this, OpenAI introduced a “Feedback Contest” encouraging users to provide feedback on problematic model outputs during their interactions with ChatGPT. Many corporations with generative AI products have also turned to “red-teaming” events, where users rigorously test models to identify vulnerabilities and harmful biases (such as the Red Team village at DEFCON, a hacking conference) (CON, 2024). In response to calls from government agencies like the Executive Order on Safe, Secure, and Trustworthy Development and Use of AI (The White House, 2023), many civil society organizations has called for technology companies to engage external domain experts, including end users, in auditing their generative AI products (Groves et al., 2023; Bogen and Winecoff, [n. d.]; Park, [n. d.]; Feffer et al., 2024).
@@
A small but growing body of research has begun exploring the needs and challenges faced by AI practitioners in conducting user-engaged AI audits. For example, through interviews with 35 AI audit practitioners, Ojewale et al. identified the need for tools to support AI practitioners in engaging those directly impacted by AI deployments, offering insights into how to better design participatory AI tools and processes with accountability (Ojewale et al., 2024). In a study with industry AI practitioners who are engaging users in auditing their AI systems, Deng et al. found that practitioners often repurpose crowdsourcing platforms for user audits but struggle to scaffold users toward productive auditing strategies and to derive actionable insights from users’ audit outcomes (Deng et al., 2023a). Many other studies on responsible AI have highlighted the lack of efficient tooling and collaboration processes for practitioners to engage end users in testing and auditing AI products and services (Madaio et al., 2020, 2024; Wang et al., 2023; Deng et al., 2022, 2023b).
@@ 3.1.1. Think-aloud study with end users
To understand the challenges end users encounter when participating in GenAI audits, and to identify opportunities to better support them beyond existing approaches (Section 2), we first conducted think-alouds with 11 end users (U01 - U11). For these formative studies, we aimed to recruit participants spanning diverse demographics and levels of technical literacy. To this end, we recruited participants through Craigslist and Nextdoor. 10 of our 11 participants had not previously used any T2I systems; participants ranged from 18 years old to 80 years old, with 47.45 average. 2 participants has high school degrees, 2 Associate degrees, 5 bachelor degrees, 1 master degree, and 1 Ph.D. We include detailed participants descriptions in Table 2 under the Appendix A.1.
@@
After this initial phase, for the next 20 minutes participants continued their testing, but now with access to a set of three low-fidelity prototypes intended to support their process (see Figure 11). The first prototype was a pairwise comparison feature, simulated by placing two browser windows side-by-side. We hypothesized that viewing prompt outputs side-by-side would help participants, especially those unfamiliar with T2I systems, understand how prompt changes affect the output. The second prototype was a sidebar displaying a sample user audit report in text format. This report featured an election-themed prompt and image pair, with a brief sentence explaining why the AI-generated image could be harmful. We hypothesized that this example could inspire participants to search for similar cases and better articulate the harms they identified. The third prototype was a T2I harm taxonomy, mapping Shelby et al.’s sociotechnical AI harm taxonomy (Shelby et al., 2023) to concrete examples of harmful T2I behaviors, based on examples from (Luccioni et al., 2023). We hypothesized that this taxonomy would help users brainstorm a broader range of potential harms to various communities and use terminology to describe these harms effectively.
@@ 3.1.2. Semi-structured Interview with Industry GenAI Practitioners
To better understand industry GenAI developers’ needs for support in effectively engaging end-users in AI auditing, we conducted semi-structured interviews with 7 industry AI developers (P01 - P07). To recruit participants, we adopted a purposive sampling approach (Campbell et al., 2020), with the aim of recruiting industry practitioners who had already attempted to engage end users in evaluating and auditing their GenAI products  [^2] Practitioners frame this activity differently depending on their organization, in some cases using terminology like “red-teaming”.. We recruited participants through social media (e.g., LinkedIn and X), and via direct contacts at large technology companies. Overall, industry practitioners come from diverse backgrounds, working in technical, user-facing, and managerial roles on various types of AI products powered by Large Language Models (LLMs) or other types of generative AI. Notably, all of them are currently using crowdsourcing platforms to engage individuals outside their AI teams in auditing or red-teaming their AI products and services. Table 4 in the Appendix A.1 provides an overview of participants’ backgrounds  [^3] Following prior work studying responsible AI practices in industry (Madaio et al., 2020; Rakova et al., 2021), to protect participants’ confidentiality, we omit potentially identifying demographic details such as gender and age, and abstract our descriptions of participants’ companies and roles..
@@ 4.1. The WeAudit Workflow
As shown in Figure 4, the WeAudit workflow is organized into two intersecting, iterative loops: Investigate and Deliberate. The overall loop structure of this workflow is informed by existing models of information foraging and sensemaking (Pirolli and Card, 1999), and these loops are further broken down into six sub-activities: Explore, Inspect, Reflect, Report, Discuss, and Verify. In particular, the idea of the loop draws inspiration from Pirolli and Card’s model of how intelligence analysts gather evidence, generate hypotheses, re-evaluate insights, and seek additional evidence in an iterative manner(Pirolli and Card, 1999). We conceptualize AI auditing as involving a similarly structured process in which auditors systematically inspect, organize, refine, and validate their insights (Pirolli and Card, 1999; Cabrera et al., 2022).
In the Investigate loop, user auditors may begin investigating AI harms by exploring prompts that could lead to harmful outputs (DG1, DG3). Auditors inspect the resulting AI-generated outputs (DG1), and reflect on potential biases or harmful impacts, drawing upon their unique knowledge and experiences (DG2). Auditors can then report their findings, documenting potential harms they identified (DG4). In the Deliberate loop, user auditors can discuss their findings with other auditors or the broader public (DG5). This collaboration allows for a more holistic understanding of AI-related harms, leveraging user auditors’ diverse perspectives. Additionally, reported findings can be verified by other auditors or the public (DG6), to provide a sense of whether a given observation is perceived as potential harmful by multiple people.
@@
In the workflow, we use the circular arrows to highlight that the entire workflow is iterative, following prior sensemaking research (Pirolli and Card, 1999). For example, new insights that arise during discussion or verification may inspire user auditors to conduct new investigations. Moreover, this iterative workflow engages user auditors in fluidly moving back-and-forth between individual and collective activities, where individual efforts contribute to the collective understanding of AI behaviors and vice versa. Throughout the entire process, user auditors are encouraged to reflect on potential harms based on their own lived experiences and identities. Inspired by past work on collaborative sensemaking (kriplean2012supporting), the “reflect” and “report” activities function as bridges between Investigate and Deliberate loops. User auditors’ reflections and reports on potential harms provide material for collective discussion and deliberation. In turn, discussions around these reflections and reports may inspire further investigations. Given this bridging role, we represent “reflect” and “report” at the intersection of these loops.
@@ 4.2. The WeAudit System
Figure 5. WeAudit interface for features: (a) Pairwise Comparison (Section 4.2.1), (b) Prompt History Sidebar (Section 4.2.2), (c) Worked Examples Repository (Section 4.2.3), and (d) Social Augmentation (Section 4.2.4)
@@ 4.2.4. Social Augmentation
In the formative study, we found initial evidence that participants found the “sidebar examples” from other user auditors useful (see Section 3). Prior research in CSCW, crowdsourcing, and sensemaking also shows that reviewing other users’ behaviors can affect users motivations and strategies for completing a task (Morris and Horvitz, 2007; Amershi and Morris, 2008; Dow et al., 2012). Inspired by this, we provide users with an aggregated view of what other users have been reporting. User auditors can click on the ”What are other users auditing?” button to view a distribution of “affected groups” (e.g., religion or sexual orientation) and “types of harms” (e.g., stereotyping social groups or economic loss) currently covered by submitted audit reports, with ones that are “underexplored” and ones that are “most explored” highlighted (Figure 5). By clicking each tag, users can view the specific user audit posts associated with those tags. Through this form of social augmentation, our goal is to encourage users to explore auditing directions that have been less frequently explored by others (DG3), drawing upon their own unique identities, experiences, and knowledge (DG2).
Figure 6. WeAudit Interface for features: (e) Audit Report Portal (Section 4.2.5), an example of audit report authored by a user auditor in our user study through the Audit Report Portal, and (f) Audit Discussion Forum (Section 4.2.6). Please refer to Section 4.2.5 to review the concrete questions in the Audit Report Portal.